Skip to main content

AAVSOnet news 2014-01-11

HQA's picture
HQA
Offline
Joined: 2010-05-10

For those receiving data, please check your images.  If the exposures need to be adjusted, let me know.  This is particularly important for those stars with large amplitude, as correct exposures near minimum might saturate near maximum for example.

The major recent action has been behind the scenes, renaming files and directories, improving flat collection and creation, making the pipeline consistent between telescopes.  We've also added a few new objects and pared down other projects as requested by the researcher.  V1369 Cen is being monitored by both BSM South and APASS south, the latter obtaining a set of BVgrizY images per night when the nova is fainter than 5th magnitude.

BSM_Berry is operational, though I'm currently having an issue with synchronizing with the mount.  This should be taken care of on the next night of operation.  We're currently using an ST-10 donated by Bart Staels while the ST-8 is being repaired at SBIG.

BSM_NM is operational.  I updated FocusMax, but need to reconfigure it slightly to get it to work properly.  Until then, the system is focused at the beginning of a night and about 2hrs before dawn, so late night may show worse focus.

BSM_Hamren also has a focus problem, but this one is likely mechanical.  We hope to do some tests the next time Bob Stine is in residence.  This system will be moved in the near future to another site.

BSM_South is working fine.  Peter has had a string of good nights (dry is not always good in Australia!).

Wright30 is no longer with us.  The mount was having communication problems. Phil Coker was kind enough to donate Coker30, an upgraded LX200 system, installed at Bill Goff's observatory in place of W30.  We are just about finished with commissioning this new system, and it will be taking more consistent data now that the holidays are over.

Bareket has supplied a few V339 Del images, but their photometric filters have degraded and the data quality is not high enough for other projects.  We've ordered replacement filters and we'll upgrade Bareket in a few weeks.

SRO is fully operational.

OC61 is fully operational.  They are progressing with their instrument selector, so we are preparing the eShel spectrograph for initial tests in the near future.

TMO61 is operational, but is not run when Jon Holtzman is out of town.  We're debugging a new Boltwood cloud sensor so that the system can be used on more nights.  The RA drive does not work as reliably as we would like, so that is being debugged.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484